# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

### v1

@@ 3.3 Continual Learning Algorithms
We benchmark the continual learning baselines below (Fig.4b). Naive (lower bound) with the SGTR model as the backbone, is trained on each task in sequence without any measures to prevent catastrophic forgetting. EWC(Kirkpatrick et al., 2017) is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks. PackNet(Mallya & Lazebnik, 2018) is a parameter-isolation method, where it iteratively prunes and pre-trains the network parameters so that it can sequentially pack multiple tasks within one network. Replay(Rolnick et al., 2019) includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario. We vary $M=$ 10%, 20%, and 100%. We randomly select images present in the current task and add them to the buffer. As the image may contain multiple object/relationship ground truths, the number of replays on ground truth notations in each task might vary. See Fig. S1(c)(d), Fig. S2(c)(d) and Fig. S3(c)(d) in Sec. A.3.4 where we report the number of ground truth notations stored in the memory buffer for each task in each learning scenario. However, these variations do not interfere with fair comparisons across replay methods with different memory buffer capacities. We also introduce Joint Training — an upper bound where the SGG model is trained on the entire CSEGG dataset.
To alleviate the problem of long-tailed data distribution, we also introduce BLS and LVIS in replay methods (Replay + BLS and Replay + LVIS). In the data re-balancing replay, the data in the memory buffer of Replay gets first re-sampled and then mixed with the data from the current task for training. As LVIS is an image-level up-sampling technique, the images containing the tail classes in the memory buffer are over-sampled. In contrast, BLS adds an extra instance-level under-sampling technique, which under-samples instances of head classes on the replay images. With these two approaches, the exact number of ground truth notations per image in the memory buffer varies over tasks in each learning scenario. We report these numbers in the Fig. S4 in Sec. A.3. The total number of instance replays for Replay+LVIS is slightly larger than the ones for Replay+BLS. We show the results of this effect in Sec. 4.2.
@@ 3.4 Evaluation Metrics
Same as existing SGG works (Xu et al., 2017; Li et al., 2022b), we adopt the evaluation metric recall@K (R@K) on the predicted scene graphs $G$. As CSEGG is long-tailed, we further report the results in mean recall (mR@K) over head, body, and tail relationship classes in Sec. A.4.2. For consistency, we provide the R@K results in the main text, while we provide and analyze results in mR@K in Fig. S8 and Fig. S9 in Sec. A.4.2.
Figure 5: Results of average recall and forgetting over tasks in Learning Scenarios 1 (a and b) and 2 (c and d) based on the SGTR backbone as the SGG model. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@20, the better. See Fig. S5,S6 in Sec. A.4.1 for results of Avg.R@K, F@K, and mR@K, when K$=50,100$. See Sec. A.4.3 for results on FWT@K and BWT@K. See Sec. A.5 for the results based on the CNN-SGG backbone (Sec.3.2).
To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1. An ideal CSEGG model could maintain the same $R@K$ on $D_{t=1}$ over tasks; thus, $F=0$ for all tasks. The more negative F is, the more severe in forgetting an model gets. To assess the overall recall of CSEGG models over tasks, we also report the continual average recall (Avg. R@K). Avg. R@K is computed as the average recall on all the data at the previous and current tasks $D_{i}$, where $i\in\{1,2,...,t\}$.
To assess whether the knowledge at previous tasks facilitates learning the new task and whether the knowledge at new tasks enhances the performances at older tasks, we introduce Forward Transfer (FWT@K) (Lin et al., 2022) and Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017). See Sec. A.2.3 and Sec. A.4.3 for its definitions, results, and analysis.
@@
4 Results
4.1 Continual Scene Graph Generation Remains a Great Challenge.
We present Avg. R@20, F@20, FWT@20, and BWT@20 results for learning scenario 1 (S1) in Fig. 5 (a)(b) and Fig.S11(a)(b). Notably, all continual learning baselines start from a similar Avg.R@20 in Task 1 and their performance drops over subsequent tasks. This implies that catastrophic forgetting about learned relationships occurs when the CSEGG models learn new relationships. Among all the methods, the naive method takes no measures of preventing catastrophic forgetting, resulting in the largest drop in Avg.R@20 and F@20. In contrast, a replay method with all the old data to rehearse in the current task (Replay(100%)) yields the least forgetting and maintains a high Avg.R@20. Surprisingly, even though Replay(100%), as an upper bound, replays all the data in the current and previous tasks, there is still a drop in performance. This could possibly be due to the long-tailed data distribution in the memory buffer, which makes the rehearsal of tail classes even less frequent in new tasks, and thus, deteriorate the recall performances of tail classes. We also compared EWC versus the replay methods. Though EWC outperforms the naive baseline in earlier tasks, it fails in longer task sequences. Different from EWC, Replay with 10% still achieves a higher Avg.R@20 score of 8.55% and a higher F@20 of -22.21%. This aligns with the existing continual learning literature that replay methods are more effective than weight-regularization methods in eliminating catastrophic forgetting (Lesort et al., 2019). PackNet is a parameter isolation method. While PackNet outperforms EWC, its performance is inferior to that of Replay(10%). As expected, we also compare the replay methods with different memory buffer sizes. Replaying more old data helps CSEGG performances. Joint training demonstrates superior performance over all the tasks in Learning Scenario 1 as seen in Fig 5. This aligns with the existing continual learning literature that joint training is a superior upper bound than Replay(100%). As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances. Aligning with the existing literature (Singh et al., 2022), we found a prominent effect of task sequences in CSEGG (Fig. S13 and Sec. A.4.4).
Learning scenario 2 (S2) approximates the real-world CSEGG setting where there are constantly new demands in detecting new objects and new relationships simultaneously. The results of S2 in Avg.R@20 and F@20 are provided in Fig. 5(c)(d). Compared with S1, the overall Avg.R@20 and F@20 drop more significantly over tasks. For example, even with 20% memory buffer size, the replay method only achieves Avg. R@20 score of 6.57% and F@20 of -17.17% in Task 2. This suggests that the real-world CSEGG remains a challenging task and there still exists a large performance gap for state-of-the-art CSEGG methods. Moreover, we also made an interesting observation that Replay (100%) outperforms the upper bound of the joint training in the first task of Scenario 2. This performance difference could be attributed to the presence of long-tailed data distribution across tasks, with the first task containing more tailed classes than head classes. This is in contrast to the task splits in Scenario 1 where both head and tail classes are uniformly sampled for every task. Consequently, joint training struggles in the first task due to sub-optimal performance in tailed classes. To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig. S18 and Sec. A.6.1) and Scenario 2 (Fig. S19 and Sec. A.6.2).
Model Avg. R@20 $\uparrow$ F@20 $\uparrow$ T1 T5 T1 T5 Replay(10%) 28.7 8.55 0 -22.21 Replay(10%)+LVIS 28.7 14.38 0 -15.39 Replay(10%)+BLS 28.7 9.56 0 -22.4 Replay(100%) 28.7 16.17 0 -12.24 Table 1: Results at Task 1 and 5 in Learning Scenario 1 when sampling techniques on long-tailed distribution are applied. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. The best results are in bold.
[Uncaptioned image] Figure 6: Generalization results in Learning Scenario 3. See Sec. 3.4 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning baselines. Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
4.2 Accounting for Long-tailed Distributions Enhances CSEGG
Due to the imbalanced data distribution in the real world, long-tailed distribution remains a unique challenge for CSEGG. Here, we introduce two data sampling techniques (LVIS and BLS) to counter-balance the long-tailed data distributions in the memory buffers as well as the feed-forward training tasks (Sec. 3.2 and Sec. 3.3). We report the results of Replay(10%)+BLS and Replay(10%)+LVIS in learning scenario S1. In Tab. 1, both long-tailed methods with replay(10%) outperform naive replays by an average margin of 3.42% in Avg.R@20 and 3.35% in F@20. This implies that data re-sampling techniques enhance general continual learning performances in long-tailed incremental learning settings. Indeed, we made the same observations after splitting the classes from each task into tail, body, and head classes and reporting their mR@K in the Fig. S10 in A.4.2. Interestingly, we see that Replay(10%)+BLS underperforms Replay(10%)+LVIS by 4.32% in Avg.R@20 and 6.42% in F@20. This contradicts the findings that BLS is more effective than LVIS in the classical SGG problem (Li et al., 2021). The performance discrepancy could be due to the difference in the number of replay instances in both approaches after these two data re-sampling methods are applied to the memory buffer (see Sec. 3.3). This emphasizes that the long-tailed learning methods explored in the SGG problem may not be effective in CSEGG. We need to explore new long-tailed learning methods specifically for CSEGG.
4.3 CSEGG Improves Generalization in Unknown Scene Understanding
@@
Notably, we also found that the SGTR model is very good at generalizing to classify relationships Fig. 6 (b). Even in Task 1, both the naive method and the Replay(10%) achieve 45% recall of known relationships among unknown objects in the generalization test set. As the CSEGG models continuously learn to detect more new objects and classify their relationships in subsequent tasks, their relationship generalization ability among unknown objects saturates around Task 3. See Fig. S20 in Sec. A.6.3 for visualization examples.

### v3

@@ 3.2 Competitive CSEGG Baselines
Continual Learning Methods. We include the following continual learning methods (Fig. S2(b)): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting. (2) EWC[22] is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks. (3) PackNet[38] is a parameter-isolation method, iteratively pruning the network parameters after every task, so that it can sequentially pack multiple tasks within one network. (4) Replay@M[48] includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario. We vary $M=$ 10%, 20%, and 100%. (5) Joint Training is an upper bound where the SGG model is trained on the entire CSEGG dataset.
Sampling Methods to Handle Long-Tailed Distribution. We adopt the five data sampling techniques to alleviate the problem of imbalanced data distribution during training. (1) LVIS[18] is an image-level over-sampling strategy for the tailed classes. (2) Bi-level sampling (BLS) [29] balances the trade-off between image-level oversampling for the tailed classes and instance-level under-sampling for the head classes. (3) Equalized Focal Loss (EFL) [27] is an effective loss function, re-balancing the loss contribution of head and tail classes according to their imbalanced distribution. EFL is enabled all the time for all the CSEGG baselines. In addition to applying data sampling techniques to the training sets, we can also apply LVIS and BLS techniques to the data stored in the replay buffer. We name these data sampling techniques applied during both training and replays as (4) LVIS@Replay and (5) BLS@Replay.
@@ 3.3 Evaluation Metrics
Same as existing SGG works [63, 28], we adopt the evaluation metric recall@K (R@K) on the top $K$ predicted triplets in the scene graphs $G$. As CSEGG is long-tailed, we further report the results in mean recall (mR@K) over the head, body, and tail classes. Forgetfullness (F), Average (Avg.) performance, Forward Transfer (FWT) [34] and Backward Transfer (BWT) [35] are standard evaluation metrics used for continual learning in image recognition and object detection tasks. In Scenario 1 and 2, we adapt these metrics to recalls R@K and introduce F@K, Avg.R@K, FWT@K, and BWT@K respectively for CSEGG settings. Similarly, we also adapt these metrics to mR@K. We explored CSEGG with K=20, 50, and 100. Since our results are consistent among Ks, we omit “@K" and analyze all the results based on K=20 in the entire text.
In scenario S3, we evaluate all CSEGG methods in the standalone generalization test set, shared over all the tasks. To benchmark generalization abilities in unknown object localization and relationship classification among these unknown objects, we introduce two evaluation metrics: Gen Rbbox@K and Gen R@K. As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted bounding box locations with Gen Rbbox@K. To evaluate whether the predicted bounding box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth. Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct. We vary IoU thresholds from 0.3, 0.5, to 0.7. To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes. See Sec. 0.A.3 for detailed definitions of these metrics. All experimental results are averaged over three runs.
4 Replays via Analysis by Synthesis (RAS)
Figure 3: Schematic of our proposed Replays via Analysis by Synthesis (RAS) method. At task $t+1$, our RAS stores all the triplet labels $U_{t}$, such as <man, on, horse>, from the previous tasks. It then re-composes these triplet labels to create in-context prompts, utilizing them as inputs to generative image models to synthesize images for replays. For predicting scene graphs on these synthesized images, we employ the frozen model $M_{t}$ from the preceding task $t$, marked with “snowflakes". Subsequently, these predicted scene graph notations, along with their corresponding synthesized images, contribute to “pseudo" replays, preventing the current model $M_{t+1}$ from experiencing forgetting. See Sec. 4 for more details.
To address the complexities in CSEGG, we introduce our “Replays via Analysis by Synthesis" method, dubbed RAS. Our RAS belongs to the group of generative replay methods for continual learning. However, different from the existing generative replay methods [52, 14, 15], our RAS leverages symbolic replays with state-of-the-art diffusion models. Moreover, rather than generating any random images for replays, our RAS is capable of generating in-context images following semantic rules, such as object co-occurrences. The schematic of our RAS is presented in Fig. 3. Next, we focus on how RAS generates images and predicts SGG annotations on these images for replays.
Image Generation. At the current task $t+1$, our RAS requires storing the frozen old model snapshot $M_{t}$ at the end of the previous task $t$ and all the triplet labels $U_{t}$, which are parsed from all the scene graphs aggregated from all the previous tasks. These triplets contain object labels, subject labels, and relationships among them. For example, <man, on, horse> and <man, in front of, horse> are two unique triplet labels. Unlike the traditional replay methods in continual learning literature, our method refrains from storing original images $I_{i}$ or scene graphs $G_{i}$ in the training sets, thereby eliminating storage issues and privacy concerns.
To generate images $I^{\prime}_{j}$ for replays, our RAS feeds text prompts, which are formed by a set of chosen triplet labels and describe the diverse in-context scenes, into the state-of-the-art Stable Diffusion model [49]. The small object on the table is most likely to be a spoon rather than an elephant. Context plays important roles in visual perceptions [5]. To generate text prompts describing context-congruent scenes, we employ a context checker. First, the context checker uses the pre-trained large language model BERT [13] to extract embeddings for each triplet label in $U_{t}$. As BERT has been pre-trained on a large corpus of text data, it learns to capture context-relevant representations of words. Next, hierarchical clustering is performed on these embeddings using the agglomerative clustering algorithm [40]. This ensures that each cluster contains only embeddings that are semantically close. The threshold for the agglomerative clustering algorithm is set to 0.6. As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for the image generation model.
In practical applications, conducting agglomerative clustering on all triplet labels in $U_{t}$ is computationally demanding, as it requires computing pairwise embedding similarities among all the triplet labels. To address this, RAS opts for a more efficient approach during replays by selecting a subset of triplet labels and clustering their embeddings. Recognizing that real-world scenes often exhibit a long-tailed distribution with certain objects or relationships being more prevalent, we introduce the Long-Tailed Distribution (LTD) module for balancing this distribution in $U_{t}$. Unlike image-level sampling methods like BLS and LVIS [29, 18] discussed in Sec. 3.2, our LTD module in RAS operates at the triplet level. For each triplet label, its dropout rate is determined proportionally to its frequency in $U_{t}$. Specifically, we define the drop-out rate $d_{k}$ for the $k$-th triplet as: $d_{k}=f_{k}/(\sum_{i=1}^{i=N}f_{i})*\alpha$, where $N$ is the total number of triplets in $U_{t}$, $\alpha=0.7$ is a scaling factor, and $f_{i}$ is the frequency of the $i$-th triplet in $U_{t}$. This sampling formula enables RAS to select triplets from tail classes more frequently compared to those from head classes.
To generate a text prompt from the chosen triplet labels, we employ a straightforward English language construct using the conjunction "and". This involves combining all the selected triplet labels into a sentence by starting with "Realistic Image of". For instance, if the triplet labels are <man, on, horse>, <house, behind, horse>, and <man, in front, house>, the generated prompt becomes “Realistic Image of man on horse and house behind horse and man in front of house." To increase exemplar diversity for replays, we use the Stable Diffusion model [49] to generate $\gamma$ number of images for the same text prompt. In practice, we set $\gamma=10$ over all the learning scenarios.
Figure 4: Visualizations of example generated images given the text prompt from our RAS. Each row presents two examples of generated images. Given the input text prompt on the left, the generated image is displayed on the right. These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
We provide the visualization examples of some synthesized images along with the corresponding text prompts in Fig. 4. From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence, thanks to the Stable Diffusion Model [49].
Scene Graph Prediction on Synthesized Images. During replays, to train the model $M_{t+1}$ on $I^{\prime}_{j}$, we also need to predict their corresponding scene graph notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. As the frozen model snapshot $M_{t}$ at the end of task $t$ carries prior knowledge for SGG from the previous tasks, we use it to predict notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. These $G^{\prime}_{j}$ comprises object nodes $O^{\prime}_{j}$ with their respective classes $c^{\prime}_{j}$, along with object locations $b^{\prime}_{j}$. Additionally, it includes corresponding relationship nodes $R^{\prime}_{j}$ formed by triplets <$o^{\prime}_{s}$, $p^{\prime}_{k}$, $o^{\prime}_{j}$> representing subject, predicate, and object nodes, respectively. These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars, used for replays.
5 Results
SGTR[28]
Methods Learning Scenario 1 (S1) Learning Scenario 2 (S2)
@@ 5 Results
Joint 20.15 0 4.6 0 - - 12.64 0 9.84 0 - -
Replay@100% 16.17 -12.24 3.32 -1.34 -1.77 -11.72 4.56 -4.13 4.56 -5.61 -1.045 -30.25
Naive 1.33 -28.7 0.86 -1.74 -2.03 -60.67 0.51 -23.22 0.05 -11.31 -3.77 -62.34
EWC[22] 1.89 -28.4 0.96 -1.72 -1.17 -52.45 0 -23.22 0 -11.31 -2.65 -50.12
PackNet[38] 7.19 -25.67 1.35 -1.64 -1.03 -42.35 1.67 -22.77 0.9 -10.33 -1.4 -42.45
Replay@10% 8.55 -22.21 4.33 -1.44 4.29 -38.35 1.81 -20.72 1.15 -9.64 -0.9 -40.67
Replay@20% 9.25 -20.35 4.78 -1.42 3.21 -  31.98 2.57 -17.17 1.56 -8.07 -0.67 -38.27
Ours* 10.78 -18.92 5.6 -1.39 2.3 -25.56 3.45 -10.23 2.75 -6.57 -0.54 -35.67
TCNN[63]
Methods Learning Scenario 1 (S1) Learning Scenario 2 (S2)
@@
Joint 19.53 0 3.9 0 - - 4.3 0 3.7 0 - -
Replay@100% 13.45 -8.83 3.6 -0.35 -1.5 -10.45 12.45 -4.13 3.2 -0.56 -2.1 -20.34
Naive 0.98 -21.2 0.74 -1.35 -3.45 -43.87 0 -18.22 0.45 -2.67 -4.12 -53.12
EWC[22] 2.36 -21.05 0.67 -1.34 -2.34 -39.89 0 -18.22 0.03 0 -3.77 -51.67
PackNet[38] 3.2 -19.7 1.1 -1.13 -1.3 -32.45 1.1 -17.82 0.84 -1.97 -2.84 -40.34
Replay@10% 5.67 -18.9 3.21 -1.05 1.45 -28.34 1.81 -16.72 1.03 -1.74 -1.4 -43.56
Replay@20% 6.23 -17.45 3.5 -1.01 1.01 -24.32 2.37 -15.17 1.45 -1.53 -1.1 -38.56
Ours* 7.8 -15.67 3.9 -0.95 0.5 -19.83 4.67 -11.31 2.2 -0.89 -0.97 -29.65
Table 2: Results of CSEGG for various continual learning methods applied on the two SGG backbones (SGTR and TCNN) in Learning Scenarios 1 and 2. See Sec. 3.2 for continual learning baselines. See Sec. 3.3 for evaluation metrics. The higher the evaluation metrics, the better. The best are in bold.
5.1 RAS outperforms all the CSEGG baselines in Scenarios 1 and 2
[…]
LVIS@Replay@10% 3.98 -1.54
BLS@Replay@10% 4.34 -1.47
RAS (ours) 5.6 -1.39
Table 3: Results at Task 5 in Learning Scenario 1 when sampling techniques are applied to long-tailed distribution data. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. We copy the results of our RAS from Tab. 2 for easy comparisons. The best results are in bold.
Figure 5: Results in Learning Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
The results for Avg. R, F, mR, mF, FWT, and BWT in learning scenarios 1 (S1) and 2 (S2) are presented in Tab. 2. Our observations align with established research in continual learning, especially in image classification and object detection: regardless of the SGG architectures, over both learning scenarios, Naive consistently performs the worst, showcasing significant catastrophic forgetting. Rehearsal-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet. However, none of them surpass our RAS.
Moreover, compared to Replay@20%, RAS achieves superior performance while requiring a smaller constant storage size, equivalent to storing 15% exemplary images from the training set in the Replay methods. This storage size is determined by the image dimensions of 800 by 600 pixels and the number of model parameters in RAS. It is also worth noting that RAS only requires storing the old model snapshot, the triplet labels in $U_{t}$, and the image generative model, eliminating privacy concerns associated with storing original images.
In Learning Scenario 2 (S2), the task involves classifying both new objects and new relationships, significantly escalating the level of difficulty compared to S1. As evident from Tab. 2, all CSEGG models, including Replay@100%, exhibit a decline in overall performance when compared to the upper bound Joint. Although our RAS achieves the leading performance among all the baselines, its performance is still far from Joint. This underscores the persistent challenge posed by S2 in the context of CSEGG. Future work should explore new approaches to address the complexities inherent in this scenario.
To tackle the issue of imbalanced data distribution in real-world scenarios, we incorporate two established data sampling techniques (LVIS@Replay@10% and BLS@Replay@10%) into our experiment (Sec. 3.2). The outcomes in learning scenario S1 are presented in Tab. 3. We observed that their performance falls short of our RAS, underscoring the efficacy of RAS in addressing long-tailed distribution during generative replays. We also noticed that BLS@Replay@10% significantly outperforms LVIS@Replay@10%, contrary to findings in the classical SGG problem where BLS is considered more effective than LVIS [29]. The performance difference may stem from variations in the number of replay instances between the two approaches after applying these data re-sampling methods to exemplar images in the memory buffer (Sec. 3.2). This observation suggests that the original sampling methods designed for addressing long-tailed distributions in the classical SGG problem may not be as effective when applied to CSEGG.
As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances. Aligning with the existing literature [54], we found a prominent effect of task sequences in CSEGG (Fig. S3 and Sec. 0.A.4.1). To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig. S5 and Sec. 0.A.5.1) and Scenario 2 (Fig. S6 and Sec. 0.A.5.2).
5.2 CSEGG Models Can Generalize in Unknown Scenes
Fig. 5 illustrates the generalization results for detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3). In Fig. 5 (a), an increasing trend in Gen Rbbox is observed with the increasing task number for all CSEGG methods, indicating improved generalization in detecting unknown objects. Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with IoU = 0.7, showcasing the SGTR model’s ability to generalize “objectness". As expected, with an increase in IoU threshold from 0.3 to 0.7, we found that Gen Rbbox decreases due to fewer bounding boxes being considered correct. Moreover, we also compared the generalization performance in object detection between Replay@10% and Naive. Contrary to previous observations in S1 and S2 (Tab. 2), we found that Replay@10% show a decline in Gen Rbbox, possibly due to a fixed number of detected object bounding boxes output by CSEGG methods. Similarly, our RAS exhibits a reduced G Rbbox compared to Replay@10% and Naive, likely for the same underlying reason.
In Fig. 5 (b), Replay@10% outperforms Naive in Gen R when considering correctly detected unknown object locations, emphasizing that minimizing forgetting in continual learning enhances the SGTR model’s overall relationship generalization in unknown scene understanding. However, the performance of Replay@10% is still inferior to our RAS method; implying that our RAS is more proficient in generalizing to classify relationships among unknown objects. Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize “relationship". Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig.S7 and Sec.0.A.5.3.
5.3 Ablation Studies on Our RAS Reveal Key Design Insights
$\gamma$ Context LTD Triplet Avg.R $\uparrow$ F $\uparrow$ mR $\uparrow$ mF $\uparrow$
A1 10 Multiple 8.23 -21.35 3.98 -1.98
A2 10 Multiple 9.75 -20.65 4.12 -1.65
A3 10 Single 7.75 -22.45 3.12 -2.48
A4 2 Multiple 2.45 -27.43 0.45 -9.84
A5 4 Multiple 5.67 -26.42 2.41 -3.26
A6 8 Multiple 7.89 -22.42 3.89 -2.17
Ours 10 Multiple 10.78 -18.92 5.6 -1.39
Table 4: Ablation results of our RAS on learning scenario S1 reveals key design insights. This table presents the results of ablation studies conducted to identify key components of our method, as discussed in Sec. 5.3. Results in Avg.R, F, mR, mF are reported after the last task in the scenario S1.
We introduce our default method designs in Sec. 4. Here, we vary the components in our RAS to reveal key design insights.
We propose a context checker in RAS. Here, we conduct an ablation by removing this module. Triplet labels are randomly selected and combined for text prompts. In A1 of Tab. 4, we observe a performance decrease of approximately 2% across all evaluation metrics, compared with our RAS. This suggests that generating images adhering to real-world context rules is crucial for replays. The lower performance may be attributed to the challenge of generating good-quality out-of-context images for Stable Diffusion Models and the potential domain differences affecting the SGG model $M_{t}$ in predicting out-of-context SGG notations.
The LTD sampling module in our RAS is designed to balance the distribution of head and tail triplet labels from $U_{t}$. Here, we remove the LTD sampling module and report the performance of the ablated method in A2 of Tab. 4. Compared to our RAS, we observe an absolute decrease of 1-2% across all metrics. Notably, the relative decrease is more pronounced in mR and mF than Avg.R and F. As mR and mF indicate mean Recall and mean Forgetfulness over both tail and head classes, the larger drops in these metrics suggest that the absence of LTD sampling significantly hinders the SGG model’s ability to predict tail classes from previous tasks.
In our RAS, we employ multiple triplet labels to construct text prompts for image generation. In contrast to single triplet labels, our approach yields rich text descriptions of complex scenes, allowing the SGG model to capture intricate relationships among multiple objects in the same scene. Additionally, using multiple triplets is more efficient in rehearsing, as it enables the model to practice predicting multiple triplets simultaneously within the same number of synthesized images. Indeed, when we replace multiple triplet labels with single triplet labels for text prompts, we note a decrease of approximately 3% across all metrics (compare A3 with ours in Tab. 4).
Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS. We vary $\gamma$ from 2 to 8 and present their corresponding model performances in A4-6 of Tab. 4. As expected, the SGG model’s performance improves with higher $\gamma$, indicating that generating more images increases sample diversity for rehearsals and hence, it enhances CSEGG performance. We set $\gamma=10$ as the default; however, in practical scenarios with ample compute resources, dynamically synthesizing more images could further improve CSEGG performances. This also highlights the advantage of our RAS, as it can dynamically generate numerous images for replays without expanding its storage usage.